Cordova manufactures three types of stained glass window, cleverly named Products A, B, and C. Information about these products follows: Product C Sales price Variable costs per unit Fixed costs per unit Required number of labor hours Product A S46.00 22.00 8.00 1.50 Product B $56.00 12.25 8.00 2.50 $86.00 38.00 8.00 4.00 Cordova currently is limited to 50,000 labor hours per month. Cordovas marketing department has determined the following demand for its products: Product A Product B Product C 13,000 units 9,000 units 5,000 units Required: Given the companys limited resource and expected demand, compute how many units of each product Cordova should produce to maximize its profit. (Enter the products in the sequence of their preferences; the product with first preference should be entered first. Round your answer to the nearest whole number.) Product Units Produced
Answer #1
Calculations Particulars Product A Product B Product C
A Sales price         46.00         56.00         86.00
B Variable costs         22.00         12.25         38.00
C = A - B Contribution per unit         24.00         43.75         48.00
D Required number of labor hours           1.50           2.50           4.00
E = C / D Contribution per labor hour         16.00         17.50         12.00
Ranking of production                 2                 1                 3

So, as per above calculation, contribution per unit of labor hour (limited resource) is highest of Product B and then Product A and Product C

So, Product B should be manufactured first up to the maximum extent possible

Maximum production of product B = Maximum demand = 9,000 units

Labor hours required for 9,000 units

= Units x Labor hour required per unit

= 9,000 x 2.5

= 22,500 labor hours

Remaining Labor hours

= Total labor hours – Used in product B

= 50,000 – 22,500

= 27,500 labor hours

Second rank is of product A

Total maximum product of product A = Total demand = 13,000 units

Labor hours required for production of 13,000 units

= 13,000 x 1.50

= 19,500 labor hours

Remaining labor hours

= 27,500 – 19,500

= 8,000

Product C will be manufactured with remaining labor hours

So, Number of units manufactured of product C

= Remaining labor hours / Labor hours required per unit

= 8,000 / 4

= 2,000 units